Find the equation, in standard form, of the line perpendicular to 2x - 3y = -5 and passing through (3, -2). Write the equation in standard form, with all integer coefficients. Show steps
Answer by HyperBrain(694)

2x - 3y = -5
2x=3y-5
3y=2x+5
y=2/3x+5/3
Thus, the slope of this line is 2/3.
So the slope of any line perpendicular to it is -3/2.
We use here the point slope form.
Since the unknown line passes through (3,-2),
y+2=-3/2(x-3)
2y+2=-3(x-3)=-3x+9
3x+2y-7=0>>Ans.
Power up,
HyperBrain!
Stop this mess! Click here